OpenStudy (anonymous):

If p+r = 2q and (1/q)+(1/s)=(2/r), then p:q is ?

OpenStudy (anonymous):

divide first eq. by q p/q + r/q = 2 multiply 2nnd eq. by r r/q + r/s =2 r/q = 2-r/s combine p/q + (2-r/s) = 2 p/q = r/s

OpenStudy (anonymous):

Thanks henkjan :D